True or False? If the sample size is equal to 40 and the true proportion is equal to 0.25, then the z-distribution can be used instead of the binomial distribution for calculating probabilities.

True or False? When sampling from a normal distribution where the mean= u and variance= o^2, then the sampling distribution for x will also have a mean= u and a variance= o^2
Answer #1

1)

p=0.25

n=40

z-distribution can be used instead of the binomial distribution if n*p & n*q > 5

n*p = 0.25*40=10 > 5

n*q = 0.75*40= 30 > 5

Therefore z-distribution is use instead of binomial distribution

Ans:- True

2) True

If sampling from a normal distribution where the mean= u and variance= o^2, then the sampling distribution for x will also have a mean= u and a variance= o^2

Note:-

Sampling distribution of xbar is